HYDROGEN PRESSURE BROADENING OF AMMONIA INVERSION TRANSITIONS FROM 10 - 40 K

\underline{D. R. WILLEY},R. E. TIMLIN JR.,P. L. PONDILLO and M. DERAMO, Dept. of Physics, Allegheny College, Meadville, PA 16335.

We will present pressure broadening results for the (J, K) = (1, 1), (2, 2) and (3, 3) inversion transitions of NH_{3} broadened by H_{2} at temperatures of 10 - 40 K. These results were obtained using a quasi-equilibrium collisionally cooled cell. When compared to previous low temperature He pressure broadening cross sections of NH_{3}, the cross sections for low temperature H_{2} broadening are significantly larger. We have also used existing NH_{3} - H_{2} potential surfaces to calculate theoretical cross sections for comparison with the experimental data.